OpenStudy (anonymous):

Given f(x)=x+3 and g(x)=x^2, which function represents f(g(x))?

zepdrix (zepdrix):

\[\large \color{blue}{f(x)=x+3},\qquad\color{orangered}{ g(x)=x^2}\] \[\huge \color{blue}{f(}\color{orangered}{g(x)}\color{blue}{)}= \color{blue}{f(}\color{orangered}{x^2}\color{blue}{)}\] \[\huge \color{blue}{f(x)=(x)+3} \quad \rightarrow \quad\color{blue}{f(}\color{orangered}{x^2}\color{blue}{)=(\color{orangered}{x^2})+3}\]

zepdrix (zepdrix):

Understand how I plugged that in? Lemme know if you're still confused.

OpenStudy (anonymous):

I do! Thanks so much!